# Discrete Distribution Networks

### v1

@@ 3.3 Applications
To implement ZSCG, we replace the Guided Sampler in Fig. 2(a) with a Conditional Guided Sampler. For instance, when generating an image of class $y_{i}$ guided by a classifier, we replace the “L2 Distance to Training Image” in the Guided Sampler of Fig. 2(a) with the probability of each output image belonging to class $y_{i}$ according to the classifier. We then replace “argmin” with “argmax” to construct the Guided Sampler for this classifier.
For super-resolution and colorization tasks, we construct a transform that converts the generated images to the target domain (low-resolution or grayscale images). This approach significantly reduces the impact of the missing signal from the condition on the generated images, enabling DDN to perform even very low-resolution super-resolution tasks.
@@
Efficient Data Compression Capability. The latent of DDN is a highly compressed discrete representation, where the information content of a DDN latent is $L\times\log_{2}{K}$ bits. Taking our default experimental values of $K$=512 and $L$=128 as an example, a sample can be compressed to 1152 bits, demonstrating the efficient lossy compression capability of DDN. We hypothesize this ability originates from two aspects: 1) the compact hierarchical discrete representation, and 2) the Split-and-Prune operation makes the probabilities of each node as equal as possible, thereby increasing the information entropy [26] of the entire latent distribution and more effectively utilising each bit within the latent.
In our experiments, we set $K$=512 as the default, considering the balance between generation performance and training efficiency. However, from the perspective of data compression, setting $K$ to 2 and increasing $L$ provides a better balance between representation space and compression efficiency. We refer to DDN with $K$=2 as Binary DDN (BinDDN). To the best of our knowledge, BinDDN are the first generative model capable of directly transforming data into semantically meaningful binary strings. These binary strings represent a leaf node on a balanced binary tree.
3.4 Techniques
Learning Residual. In the context of utilizing the Single Shot Generator structure, a Discrete Distribution Layer is introduced every two convolution layers. Given such small amount of computation between adjacent layers, directly regressing the images themselves with these convolutions becomes challenging. Drawing inspiration from ResNet [8], we propose a scheme for the network to learn the residual between the output images from the preceding layer and the ground truth. The output of the current layer is then computed as the sum of the output from the previous layer and the current layer’s residual. This approach alleviates the pressure of the network to represent complex data and enhances the flexibility of the network.
@@ 4 Experiments
Method Type One-Step Generation FID↓
DCGAN[21]
GAN
37.1
BigGAN[2]
GAN
14.1
IGEBM[6]
EBM
38.0
Denoising Diffusion[28]
EBM
3.2
VAE[16]
VAE
106.7
Gated PixelCNN[29]
AR
65.9
GLOW[14]
Flow
46.0
DDN(ours)
DDN
52.0
Table 1: Quantitative comparison on CIFAR-10
Figure 7: Zero-Shot Conditional Generation on CIFAR-10 Guided by a Pretrained Classification Model. Each row corresponds to a class in CIFAR-10. Specifically, the first row consists of airplanes, the third row displays birds flying against the blue sky, and the last row presents trucks. Our model successfully generates reasonable images for each class without any conditional training, demonstrating the powerful zero-shot generation capability of our DDN.
Model FID↓ rFID-FFHQ↓ rFID-CelebA↓
K=512 (default) 43.1 26.0 33.2
K=64
47.0
32.3 38.7
K=8
52.6
40.9 49.8
K=2 (BinDDN)
66.5
38.4 70.6
w/o Split-and-Prune
55.3
31.2 34.7
w/o Chain Dropout
182.3
26.5 37.4
w/o Learning Residual
56.2
40.2 40.2
w/o Leak Choice
56.0
34.3 32.2
We trained our models on a server equipped with 8 RTX2080Ti GPUs, setting the Chain Dropout probability to 0.05 by default. For the 64x64 resolution experiments, we utilized a DDN with 93M parameters, setting $K=512$ and $L=128$. In the CIFAR experiments, we employed a DDN with 74M parameters, setting $K=64$ and $L=64$. The MNIST experiments were conducted using a Recurrence Iteration Paradigm UNet model with 407K parameters, where $K=64$ and $L=10$. More extended experiments exploring the properties of DDN can be found in the Appendix.
@@ 4.1 Qualitative and Quantitative Results
Generation Quality. Fig. 5(a) and Fig. 5(b) depict random generation results of DDN on CelebA-HQ-64x64 [12] and FFHQ-64x64 [13], verifying the model’s effectiveness in modeling facial data. The generation quality on CelebA-HQ appears superior to that on FFHQ, which is also reflected in the lower FID score (35.4 VS 43.1). We surmise this disparity arises from CelebA-HQ’s relatively cleaner backgrounds and less diverse facial data compared to FFHQ.
Fig. 5(c) demonstrates the random generation results of DDN on CIFAR. Concurrently, we report the FID score of DDN on Unconditional CIFAR in Tab. 1, comparing it with classical methods of other generative models. The FID of DDN surpasses that of VAE [15] and Gated PixelCNN [30].
Training a conditional DDN is quite straightforward, it only requires the input of the condition or features of the condition into the network, and the network will automatically learn $P(X|Y)$. Fig. 5(d) and Fig. 5(e) show the generation results of the class-conditional DDN on CIFAR and MNIST, respectively. The conditional DDN displays a strong ability to learn the correct $P(X|Y)$ distribution.
Zero-Shot Conditional Generation. We trained a DDN model on the FFHQ-64x64 dataset and then tested the capability of the model for zero-shot conditional image generation using the CelebA-HQ-64x64 dataset, as shown in Fig. 2. We presented the experiments of text-guided ZSCG using CLIP [22] in the Appendix. Our generation process does not require gradient derivation or numerical optimization, nor does it need iterative steps.
In addition, we employed an off-the-shelf CIFAR classifier [20] to guide the generation of specific category images by a DDN model trained unconditionally on CIFAR. We want to emphasize that the classifier is an open-source, pre-trained ResNet18 model, with no additional modifications or retraining. Fig. 7 displays images of various CIFAR categories generated by the model under the guidance of the classifier.
@@ 5 Conclusion
In this paper, we have introduced Discrete Distribution Networks, a generative model that approximates the distribution of training data using an multitude of discrete sample points. DDN exhibit two intriguing properties: they support more general zero-shot conditional generation and offer highly compressed representation. Additionally, we proposed the Split-and-Prune optimization algorithm and several effective techniques for training DDN.
Acknowledgements. Special thanks to Haikun Zhang for her comprehensive support, including but not limited to paper formatting, CIFAR classification experiments. Thanks to Yihan Xie for her significant assistance in illustrations and CLIP experiments. Gratitude is extended to Jiajun Liang and Haoqiang Fan for helpful discussions. Thanks to Xiangyu Zhang for hosting the Generative Model Special Interest Group. Appreciation is also given to Hung-yi Lee for developing and offering his engaging machine learning course.
Supplementary Material
6 Further Demonstrations on Zero-Shot Conditional Generation
@@
Utilizing CLIP as Conditioning Guidance. As illustrated in Fig. 9, we leverage CLIP [22] along with corresponding prompts as conditioning cues. We use a Discrete Distribution Network (DDN) model, which is exclusively trained on the FFHQ dataset, to yield Zero-Shot Conditional Generation (ZSCG) results. Remarkably, DDN can generate corresponding images under the guidance of CLIP without the necessity for gradient computations and only requiring a single inference pass.
@@
Supplementary ZSCG Experimental Data. We have provided raw experimental data in the zip file of the Supplementary Material. These data have not been filtered nor have we applied any Reject Sampling strategy, hence there exist bad cases in certain ZSCG tasks such as facial recognition and CLIP-guided image editing. Notwithstanding, we emphasize that to the best of our knowledge, DDN is the first generative model capable of performing ZSCG tasks with non-image-based conditions without the need for gradients and iterations. The unprocessed data serve primarily to illustrate how DDN can effectively steer towards the correct direction under the guidance of non-image-based conditions.
7 Experimental Results on Generative Performance
Verify whether DDN can generate new images. As depicted in Figure Fig. 11, we compare the images that are closest in the training dataset, FFHQ, to those generated by our DDNs. It suggests that our DDNs can synthesize new images that, while not present in the original dataset, still conform to its target distribution.
Figure 11: Nearest neighbors of the model trained on FFHQ. The leftmost column presents images generated by the DDN. Starting from the second column, we display the images from FFHQ that are most similar to the generated images, as measured by LPIPS [35].
(a) With Split-and-Prune, Chain Dropout 0
(b) Without Split-and-Prune, Chain Dropout 0
(c) Without Split-and-Prune, Chain Dropout 0.05
Figure 12: Efficacy of Split-and-Prune and Chain Dropout on MNIST.
Figure 16: Conditional DDN performing coloring and edge-to-RGB tasks. Columns 4 and 5 display the generated results under the guidance of other images, where the produced image strives to adhere to the style of the guided image as closely as possible while ensuring compliance with the condition. The resolution of the generated images is $256\times 256$.
8 Latent Analysis
@@
Semantic Performance of Latents. We explored the semantic capabilities of DDN latents through a classification experiment on the MNIST dataset. Given the inherent tree structure of DDN’s latents, we employed a decision tree classification method, using fine-tuning data to assign class votes to nodes in the tree. For unassigned nodes, their class is inherited from the closest ancestor node with a class. We fine-tuned DDN’s latent decision tree using various numbers of labeled training set data, and the results on the test set are shown in Tab. 3. All experiments in this table were conducted with Recurrence Iteration Paradigm’s UNet, which has approximately 407k parameters. However, computational complexity increases proportionally with the number of output levels (L). These experiments substantiate that DDN’s latents encompass meaningful semantic information.
@@
9 More detailed experimental explanation
@@
10 Limitations
Limitations. There are two key limitations of Discrete Distribution Networks (DDNs):
  • Loss of high-frequency signals: The high level of data compression and the use of pixel L2 loss during optimization may result in the loss of high-frequency signals, causing the images to appear blurred. A potential improvement could be learning from VQ-GAN [5] and incorporating adversarial loss [4] to enhance the modeling of high-frequency signals.
  • Difficulty with complex data: While random choice can generate decent facial data, DDNs struggle with complex data, such as ImageNet [17]. One possible solution is to learn an additional prior model dedicated to sampling output nodes.

### v2

@@ 3.2 Applications
To implement ZSCG, we replace the Guided Sampler in fig. 3(a) with a Conditional Guided Sampler. For instance, when generating an image of class $y_{i}$ guided by a classifier $g_{cls}$, we replace the “L2 Distance to Training Image” in the Guided Sampler of fig. 3(a) with the probability of each output image belonging to class $y_{i}$ according to the classifier. We then replace “argmin” with “argmax” to construct the Guided Sampler for this classifier. Similar to eq. 1, the sampling method is as follow:
$k_{l}^{*}=\underset{k\in\{1,\dots,K\}}{\operatorname{argmax}}\;g_{cls}(f_{l}(\mathbf{x}^{*}_{l-1})[k])[y_{i}]$ (5)
After performing $L$ steps guided sampling and $L\times K$ steps of classification, the ZSCG results can be obtained without any gradient.
For super-resolution and colorization tasks, we construct a transform that converts the generated images to the target domain (low-resolution or grayscale). This approach significantly reduces the impact of the missing signal from the condition on the generated images, allowing DDN to successfully perform super-resolution tasks even when the source image has a resolution as low as $4\times 4$.
@@
Table 1: Quantitative comparison on CIFAR-10
Method Type FID↓
DCGAN GAN 37.1
IGEBM EBM 38.0
VAE VAE 106.7
Gated PixelCNN AR 65.9
GLOW Flow 46.0
DDN(ours) DDN 52.0
Efficient Data Compression Capability. The latent of DDN is a highly compressed discrete representation, where the information content of a DDN latent is $L\times\log_{2}{K}$ bits. Taking our default experimental values of $K$=512 and $L$=128 as an example, a sample can be compressed to 1152 bits, demonstrating the efficient lossy compression capability of DDN. We hypothesize this ability originates from two aspects: 1) the compact hierarchical discrete representation, and 2) the Split-and-Prune operation makes the probabilities of each node as equal as possible, thereby increasing the information entropy Shannon (1948) of the entire latent distribution and more effectively utilising each bit within the latent.
In our experiments, we set $K$=512 as the default, considering the balance between generation performance and training efficiency. However, from the perspective of data compression, setting $K$ to 2 and increasing $L$ provides a better balance between representation space and compression efficiency. We refer to DDN with $K$=2 as Binary DDN (BinDDN). To our knowledge, BinDDN are the first generative model capable of directly transforming data into semantically meaningful binary strings. These binary strings represent a leaf node on a balanced binary tree.
@@
3.3 Techniques
Learning Residual. In the context of utilizing the Single Shot Generator structure, a Discrete Distribution Layer is introduced every two convolution layers. Given such small amount of computation between adjacent layers, directly regressing the images themselves with these convolutions becomes challenging. Drawing inspiration from ResNet He et al. (2016), we propose a scheme for the network to learn the residual between the output images from the preceding layer and the ground truth. The output of the current layer is then computed as the sum of the output from the previous layer and the current layer’s residual. This approach alleviates the pressure of the network to represent complex data and enhances the flexibility of the network.
Model FID↓ FFHQ↓ CelebA↓
K=512 (default) 43.1 26.0 33.2
K=64 47.0 32.3 38.7
K=8 52.6 40.9 49.8
K=2 (BinDDN) 66.5 38.4 70.6
w/o Split-and-Prune 55.3 31.2 34.7
w/o Chain Dropout 182.3 26.5 37.4
w/o Learning Res. 56.2 40.2 40.2
w/o Leak Choice 56.0 34.3 32.2
@@ 4 Experiments
We trained our models on a server equipped with 8 RTX2080Ti GPUs, setting the Chain Dropout probability to 0.05 by default. For the 64x64 resolution experiments, we utilized a DDN with 93M parameters, setting $K=512$ and $L=128$. In the CIFAR experiments, we employed a DDN with 74M parameters, setting $K=64$ and $L=64$. The MNIST experiments were conducted using a Recurrence Iteration Paradigm UNet model with 407K parameters, where $K=64$ and $L=10$. DDN is implemented on the foundation of the EDM Karras et al. (2022) codebase, with training parameters nearly identical to EDM. Both code and model weights shall be released for reproducing. More extended experiments exploring the properties of DDN can be found in the Appendix.
@@ 4.1 Qualitative and Quantitative Results
Generation Quality. fig. 6(a) and fig. 6(b) depict random generation results of DDN on CelebA-HQ-64x64 Karras et al. (2017) and FFHQ-64x64 Karras et al. (2019), verifying the model’s effectiveness in modeling facial data. The generation quality on CelebA-HQ appears superior to that on FFHQ, which is also reflected in the lower FID score (35.4 VS 43.1). We surmise this disparity arises from CelebA-HQ’s relatively cleaner backgrounds and less diverse facial data compared to FFHQ.
We conduct a qualitative comparison with a recent work, the Idempotent Generative Network (IGN) Shocher et al. (2024), accepted by ICLR 2024. Since IGN was only experimented on the CelebA dataset and did not release its code, our comparison is limited on the CelebA dataset. As depicted in fig. 7, our DDN demonstrate better generation capabilities over IGN.
(a) DDN (Ours), FID=35.4
(b) Idempotent Generative Network (IGN), FID=39
Figure 7: Comparison of randomly generated images on CelebA-HQ 64x64. DDN produces images with clearer details and fewer artifacts compared to IGN Shocher et al. (2024).
Figure 6(c) showcases the random generation results of DDN on the CIFAR dataset. In parallel, we present the FID score of DDN on unconditional CIFAR in table 1, comparing it against classical generative models. It is worth noting that modeling CIFAR remains a challenging task, especially for new and under-explored generative models like DDN. For instance, the recent work IGN Shocher et al. (2024) did not conduct experiments on CIFAR. Instead, we wish to emphasize the unique capability of DDN in “more general zero-shot conditional generatio”.
Conditional Training. Training a conditional DDN is quite straightforward, it only requires the input of the condition or features of the condition into the network, and the network will automatically learn $P(X|Y)$. fig. 6(d) and fig. 6(e) show the generation results of the class-conditional DDN on CIFAR and MNIST, respectively. The conditional DDN displays a good ability to learn the correct $P(X|Y)$ distribution.
Figure 9: Zero-Shot Conditional Generation on CIFAR-10 Guided by a pretrained classification model without gradient. Each row corresponds to a class in CIFAR-10. Specifically, the first row consists of airplanes, the third row displays birds flying against the blue sky, and the last row presents trucks. Our model successfully generates reasonable images for each class without any conditional training, demonstrating the powerful zero-shot generation capability of our DDN.
Zero-Shot Conditional Generation. We trained a DDN model on the FFHQ-64x64 dataset and then tested the capability of the model for zero-shot conditional image generation using the CelebA-HQ-64x64 dataset, as shown in fig. 2. We presented the experiments of text-guided ZSCG using CLIP Radford et al. (2021) in the Appendix. Our generation process does not require gradient derivation or numerical optimization, nor does it need iterative steps. To the best of our knowledge, DDN is the first generative model that supports purely discriminative model as guide for zero-shot conditional generation.
In addition, we employed an off-the-shelf CIFAR classifier Phan (2021) to guide the generation of specific category images by a DDN model trained unconditionally on CIFAR. We want to emphasize that the classifier is an open-source, pre-trained ResNet18 model, with no additional modifications or retraining. fig. 9 displays images of various CIFAR categories generated by the model under the guidance of the classifier.
@@ 5 Conclusion
In this paper, we have introduced Discrete Distribution Networks, a generative model that approximates the distribution of training data using a multitude of discrete sample points. DDN exhibits unique property: more general zero-shot conditional generation. We also proposed the Split-and-Prune optimization algorithm and several effective techniques for training DDN. Additionally, we showcase the efficacy of DDN and its intriguing properties through experiments.
6 Acknowledgements
Special thanks to Haikun Zhang for her comprehensive support, including but not limited to paper formatting, CIFAR classification experiments. Thanks to Yihan Xie for her significant assistance in illustrations and CLIP experiments. Gratitude is extended to Jiajun Liang and Haoqiang Fan for helpful discussions. Thanks to Xiangyu Zhang for hosting the Generative Model Special Interest Group. Appreciation is also given to Hung-yi Lee for developing and offering his engaging machine learning course.
Appendix A Experimental Results on Generative Performance
Figure 10: Conditional DDN performing coloring and edge-to-RGB tasks. Benefiting from the reduction of the generative space by the condition, DDN is capable of generating high-quality images of $256\times 256$ resolution. Columns 4 and 5 display the generated results under the guidance of other images, where the produced image strives to adhere to the style of the guided image as closely as possible while ensuring compliance with the condition.
Verify whether DDN can generate new images. As depicted in fig. 11, we compare the images that are closest in the training dataset, FFHQ, to those generated by our DDNs. It suggests that our DDNs can synthesize new images that, while not present in the original dataset, still conform to its target distribution.
Figure 11: Nearest neighbors of the model trained on FFHQ. The leftmost column presents images generated by the DDN. Starting from the second column, we display the images from FFHQ that are most similar to the generated images, as measured by LPIPS Zhang et al. (2018).
(a) With Split-and-Prune, set Chain Dropout 0
(b) Without Split-and-Prune, set Chain Dropout 0
(c) Without Split-and-Prune, set Chain Dropout 0.05
Figure 12: Efficacy of Split-and-Prune and Chain Dropout on MNIST.
Appendix B Further Demonstrations on Zero-Shot Conditional Generation
@@
Utilizing CLIP as Conditioning Guidance. As illustrated in fig. 16, we leverage CLIP Radford et al. (2021) along with corresponding prompts as conditioning cues. We use a Discrete Distribution Network (DDN) model, which is exclusively trained on the FFHQ dataset, to yield Zero-Shot Conditional Generation (ZSCG) results. Remarkably, DDN can generate corresponding images under the guidance of CLIP without the necessity for gradient computations and only requiring a single inference pass.
@@
Appendix C Latent Analysis
@@
Semantic Performance of Latents. We explored the semantic capabilities of DDN latents through a classification experiment on the MNIST dataset. Given the inherent tree structure of DDN’s latents, we employed a decision tree classification method, using fine-tuning data to assign class votes to nodes in the tree. For unassigned nodes, their class is inherited from the closest ancestor node with a class. We fine-tuned DDN’s latent decision tree using various numbers of labeled training set data, and the results on the test set are shown in table 3. All experiments in this table were conducted with Recurrence Iteration Paradigm’s UNet, which has approximately 407k parameters. However, computational complexity increases proportionally with the number of output levels (L). These experiments substantiate that DDN’s latents encompass meaningful semantic information.
@@
Appendix D More detailed experimental explanation
@@
Appendix E Limitations
There are some key limitations of Discrete Distribution Networks (DDNs):
  • Loss of high-frequency signals: The high level of data compression and the use of pixel L2 loss during optimization may result in the loss of high-frequency signals, causing the images to appear blurred. A potential improvement could be learning from VQ-GAN Esser et al. (2021) and incorporating adversarial loss Creswell et al. (2018) to enhance the modeling of high-frequency signals.
  • Difficulty with complex data: While random choice can generate decent facial data, DDNs struggle with complex data, such as ImageNet Krizhevsky et al. (2012). One possible solution is to learn an additional prior model dedicated to sampling output nodes.
  • computational burden of zero-shot conditional generation: ZSCG requires $L\times K$ forward passes through the guided model, where $L$ is the number of layers and $K$ is the number of possible outputs per layer. When the guided model itself is computationally expensive, this results in significant computational overhead and prolonged generation time. However, since the discrimination process is parallelizable across the $K$-dimensional output space, batching techniques can be employed to mitigate latency. In addition, further research will be conducted to reduce the number of guided model calls during the ZSCG process.